Check the convergence of the sequence defined by 𝑒𝑛+1 = (1 + 1 𝑒𝑛 ) , 𝑒1 > 0. Note that this is the sequence associated with the continued fraction expansion of the Golden ratio.Β 


1
Expert's answer
2021-12-21T12:48:06-0500

"u_{n+1}=1+1\/u_n"


map "u\to 1+1/u" can be extended to a Moebius transformation of the Riemann sphere

"C\cup \{\infin\}:"

"z\to \frac{z+1}{z},T(0)=\infin,T(\infin)=1"

Its fixed points are:

"a=(1+\sqrt 5)/2,b=(1-\sqrt 5)/2"

obtained by solving the equation

"z^2-z-1=0"

We now introduce a new complex coordinate w on C, related to z via

"w=\phi(z)=\frac{z-a}{z-b}\implies z=\phi^{-1}(w)=\frac{a-bw}{1-w}"

The fixed points now are w = 0 and "w=\infin"

in terms of the new coordinate w the transformation T appears as

"\tilde{T}=\phi \circ T \circ \phi^{-1}" , then:

"\tilde{T}: w\to \frac{b}{a}w,\tilde{T}(0)=0,\tilde{T}(\infin)=\infin"


since

"\frac{b}{a}=\frac{3-\sqrt 5}{2}=-0.382"

we can infer that the fixed point 0 is attracting with basin of attraction all of C, while "\infin" is repelling. This allows to conclude that in the original setting all initial points "u_0\neq b"

lead to "\displaystyle \lim_{n\to \infin} u_n=a"


So, the sequence converges.